OpenStudy (anonymous):

If my volume is 2,000 what 2 numbers will make up the bass and height to make it equal?

OpenStudy (anonymous):

V = B* h = 2,000 Suppose B = h = a --> V = a * a = 2000 => a = √ 2000 = 20√5

OpenStudy (anonymous):

Thank